OpenStudy (anonymous):

Find a quadratic function with x-intercepts at 1 and -3 passing through the point (0,3)

OpenStudy (anonymous):

did you get help?

OpenStudy (wolfe8):

So you have what are also known as the roots of the curve: x=1 x=-3 When you rearrange those to have 0 on one side, you will get x-1=0 x+3=0 And how do you obtain the original curve function from the roots(factors)? When you get the equation, show me and we have one more thing to do.

OpenStudy (anonymous):

Would you then set up an equation and use the distributive property, so that you get y=(x-1)(x+3), which would then give you y=x^2+2x-3?

OpenStudy (wolfe8):

Correct. But then if you sketch the graph, you will have something like  The question wants a graph that goes through (0,3). What do you think we should do?